OpenStudy (anonymous):

A compact disc has a diameter of 11.8 cm. What is the surface area of the disc in square centimeters? In square meters? [Area of a circle =(π)(radius)^2]

OpenStudy (anonymous):

S = pi r^2 = pi (D/2)^2 = pi * D^2/4 = pi* 11.8^2 /4 = 109.35 cm^2